OpenStudy (anonymous):

Which of the following parabolas has a vertex of (-2,5) and opens downward?

OpenStudy (anonymous):

@jaelyn14 give us the equations

OpenStudy (anonymous):

A)y=-5(x-2)^2+5 B)y=-5(x-2)+5 C)y=5(x+2)^2 D)y=5(x-2)^2+5 @kyusakazaki

OpenStudy (anonymous):

@jaelyn14 there's somthing wrong vrify please

OpenStudy (anonymous):

@jaelyn14 the equation should be 5(x+2)^2 + 5 and it's not one of the choices given